1/2 + (2/3 + 3/4) - (4/5 x 5/6)
Answer by rfer(16322)

1/2+(2/3+3/4)-(4/5x5/6)
1/2+(8/12+9/12)-20/30
1/2+17/12-20/30
6/12+17/12-8/12=15/12=1 1/4